Use R markdown to solve the following questions.

Approximating φ The golden ratio φ is the positive solution of Simulate 1000 observations from Uniform( 0,2) distribution Then 1.5 1.0 0.5 0.0 0.5 00 1.0 0.5 1.0 .5 2.0 2x (the proportion of the values for which x2-x-1 〈 0 ) should be close to
Answer #1

Below is the R Code for approximating the golden ratio using uniform random variables:

> rm(list=ls(all=TRUE)) # To remove objects from R
> PHI=(sqrt(5)+1)/2;PHI # True value of Phi
[1] 1.618034
> N=1000 # No. of observations
> X=runif(N,0,2) # Uniform random variables
> Y=X^2-X-1 # Approximation
> (sum(Y<0)/N)*2 # Finding the proportion
[1] 1.628

We can see that the obtained value is 1.628 which is close to the true value 1.618.